Compute for the [PO4 ^3-] of a 0.025M H3PO4 solution
1
Expert's answer
2020-03-26T03:02:31-0400

The acidity constants of H3PO4 are the following:

1) H3PO4 = H+ + H2PO4- Ka1 = 7.25*10-3

2) H2PO4- = H+ + HPO42- Ka2 = 6.31*10-8

3) HPO42- = H+ + PO43- Ka3 = 3.98*10-13

The first equilibrium is predominant, as its constant is the highest. Hence, the concentration of H+ in solution is determined by reaction 1). Since H2PO4- is produced in the same quantity as H+, their concentrations can be calculated from Ka1:

[H+]2/([H3PO4] - [H+]) = Ka1

[H+]2 = 0.00725*0.025 - 0.00725*[H+]

solving the equation, we obtain [H+] = 0.0103 M, [H3PO4] = 0.025-0.0103=0.015 M

The concentration of PO43- is obtained from the product of acidity constants:

Ka1*Ka2*Ka3 = [H+]3*[PO43-]/[H3PO4]

0.0103*[PO43-]/0.015=1.82*10-22

[PO43-] = 2.73*10-18 M